OpenStudy (anonymous):

square root 8 over square root 3

OpenStudy (anonymous):

I thought it was 2 square root 2 over square root 3 but I must have done the order of operations wrong

jhonyy9 (jhonyy9):

how can you rewriting 8 using power of 2 ?

OpenStudy (anonymous):

Sorry I wrote the problem wrong. Correct now.

jhonyy9 (jhonyy9):

how is it ?

OpenStudy (anonymous):

My book says the answer is +/- 2 square root 6 over 3.

jhonyy9 (jhonyy9):

yes is right your book

jhonyy9 (jhonyy9):

because 8 = 2^2 *2 yes ?

jhonyy9 (jhonyy9):

but sqrt(2^2) = +/- 2 yes ?

OpenStudy (anonymous):

That is what I got too, once I redid it with the right number.

OpenStudy (anonymous):

I'm solving 3x squared = 8 using the square root property. Reduced it to x square = 8/3.

jhonyy9 (jhonyy9):

3x^2 =8 x^2 = 8/3 x_1,2 = +/- sqrt (8/3)

jhonyy9 (jhonyy9):

on this way ?

OpenStudy (anonymous):

Yes, that's my understanding of how to solve for x. But book says otherwise.

jhonyy9 (jhonyy9):

but sqrt8 = 2sqrt2

OpenStudy (anonymous):

That's what I thought. I'll ask instructor next class period. thanks for your help.

jhonyy9 (jhonyy9):

than will be +/- 2sqrt2 /sqrt3 = +/- 2sqrt6 / 3 so this you will get after ratzionlized the denominator ok ?

jhonyy9 (jhonyy9):

so than your book is right sure

OpenStudy (anonymous):

Oh cause you can't leave a square root on the bottom equation if I'm remembering correctly. That's a step I forgot.

jhonyy9 (jhonyy9):

yes when there is on denominator an radical you need ratzionaliz the denominator and in this way will be eliminated the radical from there ok ?

OpenStudy (anonymous):

Got it. Thanks

jhonyy9 (jhonyy9):

yw good luck bye